OpenStudy (anonymous):

Which of the following is a solution of 3x2 = 7x - 3?

OpenStudy (anonymous):

A. \[\frac{ -7\pm \sqrt{13} }{6 }\] B. \[\frac{ 7\pm \sqrt{85} }{6 }\] C. \[\frac{ -7\pm \sqrt{85} }{6 }\] D. \[\frac{ 7\pm \sqrt{13} }{6 }\]

OpenStudy (kidrah69):

We need an abc value first turn this into standard form

OpenStudy (anonymous):

I don't know how to do that. Can you help me?

OpenStudy (kidrah69):

Ok this is standard form \[ax^2+bx+c=0\]

OpenStudy (kidrah69):

here you will need to subtract 3x^2

OpenStudy (kidrah69):

what will your equation look like now?

OpenStudy (anonymous):

\[0 = 7x -3 - 3x^2\]

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

Ok so now what are your abc values?

OpenStudy (anonymous):

a=3 b=7 and c =-3

OpenStudy (kidrah69):

nope

OpenStudy (anonymous):

oops. a=-3 too

OpenStudy (kidrah69):

yep

OpenStudy (anonymous):

oh

OpenStudy (kidrah69):

relist the abc values now :)

OpenStudy (kidrah69):

\[x= \frac{ -b ± \sqrt(b^2 - 4ac) }{ 2a }\]

OpenStudy (kidrah69):

Thats the quadratic formula

OpenStudy (kidrah69):

Now all we do is plug in the values

OpenStudy (kidrah69):

Here since you may not know how... \[x= \frac{ -(7) ± \sqrt((7)^2 - 4(-3)(-3)) }{ 2(-3) }\]

OpenStudy (kidrah69):

The rest you just have to solve

OpenStudy (anonymous):

Ok. I'll try to simplify. If I have a hang up can I tag you?

OpenStudy (kidrah69):

yes :)

OpenStudy (anonymous):

Ok. So I solved that one but I'm stuck on another.

OpenStudy (anonymous):

I've gotten to the point of \[x = \frac{ -8\pm \sqrt{120} }{ 4 }\]

OpenStudy (kidrah69):

Ok wait the c value is positive

OpenStudy (kidrah69):

give me a oment i will type in the work

OpenStudy (anonymous):

But the choices are \[-2 \pm \sqrt{2}\] \[-2 \pm2\sqrt{2}\]

OpenStudy (anonymous):

and 2 others. how do i simplify further

OpenStudy (kidrah69):

http://prntscr.com/73fifz

OpenStudy (kidrah69):

get it?

OpenStudy (anonymous):

I get it but it's telling me that I need to simplify more. I don't know where to go.

OpenStudy (anonymous):

@kidrah69

OpenStudy (kidrah69):

hmmm

OpenStudy (kidrah69):

Oh you need to solve it too?

OpenStudy (anonymous):

The original was 7+2x^2 - 8x = 0

OpenStudy (anonymous):

After using the quadratic formula I got \[x = \frac{ -8\pm \sqrt{120} }{ 4 } \]

OpenStudy (kidrah69):

Why didnt you post that?

OpenStudy (kidrah69):

Whats the origional problem? :/

OpenStudy (anonymous):

One sec. Lemme start in a new question to start over and so that I can give you another best response.